OpenStudy (anonymous):

Change 4x - 3y = -15 to slope-intercept form and state the slope and y-intercept. You must show all work.

OpenStudy (anonymous):

y=mx+c m= slope c = y intercept

OpenStudy (callisto):

Or... Ax+By=C Slope = -A/B

OpenStudy (anonymous):

y = mx + b (slope intercept form) 4x - 3y = - 15 4x - 4x - 3y = - 4x - 15 (subtract 4x from both sides) -3y = - 4x - 15 (divide by -3) -3y/-3 = -4x/-3 - 15/-3 y = 4/3x + 5 m = slope = 4/3 y intercept = b = 5 I think I did this right, but you might want to get a second opinion.